The variance of Yo is given by the following formula:
ax, + b) = i=1 с. O D. n(a + b) .
Answer #1

(1)

Correct option:

A.

\frac{\sigma^{2} _{Y}}{n}

EXPLANATION:

As per Central Limit Theorem, Sampling Distribution of Sampling Mean is Normal Distribution with mean = \mu = mean of the population and variance =

\frac{\sigma ^{2}}{n}

where \sigma is the population standard deviation.

(2)

Correct option:

A.   n X a X \bar{x} +n X b

EXPLANATION:

\sum_{i=1}^{n}(ax_{i}+b)= a\sum_{i=1}^{n}x_{i}+(n\times b)

=a\times (n\times \bar{x})+(n\times b)